An adiabatic chamber has frictionless insulated piston of mass m. Mass of remaining chamber including gas is 4m. `n` moles of an ideal monoatomic gas is present inside the chamber at atmospheric pressure and temperature. Piston is in rest at equilibrium and confined to move along the length of the cylinder. A particle of mass m moving horizontally with speed v, strikes elastically with the piston. Find the change in temperature of the gas when compression of gas is maximum

Grade:11

1 Answers

saransh
7 years ago
this is a basic question 
given : adiabatic hence Q= 0 
co efficient of restitution e=1
friction = 0
v =initial vel of ball  
restitution eqn   v= u2 –  u1
momentum conservation v = u2+ u1
u1= u2= v/2  
energy conservvation  ½ m(v/2)2 = 3/2 nRT1 – T2+ 0
WD + Q = E
T1 – T= mv2/12nR